# Recovering the Pre-Fine-Tuning Weights of Generative Models

### v1

@@ 1 Introduction
To illustrate our setting, let us consider a Large Language Model (LLM). While the pre-trained version of the LLM exhibits advanced language understanding and generation capabilities, it is unaligned with human preference and is often deemed unsafe. These unsafe models can be used for example to get instructions for building a bomb or other malicious activities. To improve instruction following and enhance safety, model creators perform an alignment fine-tuning stage. Usually, only the aligned version of the LLM is published, and the recovery of the original Pre-FT unsafe weights, is implicitly assumed to be impossible. While for existing models the recovery of the Pre-FT weights poses a security and safety vulnerability; for future superhuman models, it may lead to catastrophic consequences.
@@
Figure 2: Mistral DPO Results: Our method, Spectral DeTuning, recovers the pre-fine-tuning generation capabilities with high precision, essentially undoing the DPO alignment LoRA fine-tuning. In green exact recovery, in red unrecovered words. Notice that unlike the baselines, our method recovers the semantic meaning even when the generated text is not a perfect match. Best viewed in color
Motivated by the above, we propose the task of Pre-Fine-Tuning Weight Recovery. In this paper, we tackle this task in cases where multiple LoRA fine-tuned flavors of the same source model are available. We present an overview of our setting in Fig. 1. This task is particularly timely due to two trends: i) Popular foundation models come in multiple flavors. E.g., LLaMA 2, Code LLaMA, Code LLaMA-Python, Code LLaMA-Instruct. ii) LoRA is becoming a key component for creating SoTA models [26]. These two trends have not yet merged, i.e, we are not aware of multi-flavored foundational models that use LoRA alignment fine-tuning. Here, we bring to the attention of the community the risks and perils involved in merging these trends.
@@ 4 Problem Definition
Attack setting. The attacker has access to the weights of $n$ different $\mathcal{F}_{\mathcal{P}}$ models of rank $r$, all LoRA fine-tuned from the same pre-trained source model $\mathcal{P}$. The attack succeeds with precision $\epsilon$ if the attacker can accurately recover the weights of the pre-trained source model $\mathcal{P}$ up to an $\epsilon$ precision. Formally, given $\left\{\{W_{i}^{\prime(j)}\}_{j=1}^{m}\right\}_{i=1}^{n}$, the attacker needs find $\{W^{*(j)}\}_{j=1}^{m}$ such that
@@ 5 Spectral DeTuning
We now describe our method for carrying out a Pre-FT weight recovery attack. We start by introducing our optimization objective, followed by our optimization method and finally, a rank scheduler that stabilizes the optimization and results in better convergence. For simplicity, we omit the layer index superscript $(j)$ and perform the same optimization across all layers independently.
@@ 5.2 Pre-FT Weight Recovery Algorithm
Where $\Sigma_{i|r}$ represents the top $r$ singular values $\Sigma_{i}$.
W-step. By fixing the values of $*M^{*}_{1},..,M^{*}_{n}$, we can easily compute the optimal value of $W$. The optimization problem is given by:
@@ 5.3 Rank Scheduler
The algorithm proposed in Sec. 5.2 tends to perform well in general. However, we empirically found that solving the optimization problem with high ranks can result in slow and inaccurate convergence. We therefore introduce a rank scheduler. The idea of the rank scheduler is to start by forcing $M_{i}$ to be of rank $r^{*}<r$, allowing Spectral DeTuning to focus on the most significant principal components first. $r^{*}$ is increased according to a schedule until finally $r^{*}=r$. Specifically, we use an “Increase on Plateau” type of scheduler where the rank is increased whenever the loss term from Eq. 3 plateaus. For more details see App. E. We show empirically in Sec. 7.4 that there are cases where the rank scheduler improves the rate and quality of convergence significantly.
@@ 6.1 Dataset
For each source model, we curate $15$ LoRA models fine-tuned on diverse datasets, tasks, and objectives. The dataset comprises a diverse array of layer types, including self-attention, cross-attention, and MLPs. This diversity enables us to assess the generalization capabilities of Pre-FT methods. The evaluation can be conducted on a per-model basis, per layer type, or per layer depth, allowing for a comprehensive analysis of Pre-FT methods. Overall, our dataset includes $544$ source model layers. When taking into account the fine-tuned LoRA layers, the dataset includes over $8,000$ layers. For further details see App. D.
@@ 8 Discussion and Limitations
LoRA Rank. This paper assumes a constant LoRA rank $r$ across all $n$ LoRA models, which is known to the attacker. Future work could investigate using models with varying, unknown LoRA ranks.
@@ Appendix A The Effect of the Number of LoRAs on Semantic Convergence
Figure 8: Number of LoRAs vs. Semantic Convergence - ViT
Figure 9: Number of LoRAs vs. Semantic Convergence - Stable Diffusion
Figure 10: Number of LoRAs vs. Semantic Convergence - Mistral SFT
@@
Appendix D LoWRA Bench Dataset
@@ D.2 Mistral Models
Following the original setup, use a rank $r=64$ and LoRA fine-tune the q_proj, k_proj, v_proj, and o_proj layers. This protocol results in $128$ Pre-FT model layers and a total of $128\cdot 15=1920$ LoRA fine-tuned layers for both the SFT and DPO stages. See Tabs. 7 and 7 for the fine-tuning hyper-parameters. For inference we use the following decoding hyper-parameters: max_new_tokens=50, do_sample=True, temperature=0.7, top_k=50, top_p=0.95.
@@
Table 7: Mistral DPO Hyper-parameters
D.3 Stable Diffusion Models
@@
Appendix E Spectral DeTuning Implementation Details
@@
Appendix F Runtime and Compute
@@
Appendix G Detecting the Fine-Tuned Layers
@@
Appendix H Algorithm with Rank Scheduler
@@
Appendix I Mistral Additional Results
@@
Appendix J Stable Diffusion Additional Results

### v2

@@ 1 Introduction
To illustrate our setting, let us consider a Large Language Model (LLM). While the pre-trained version of the LLM exhibits advanced language understanding and generation capabilities, it is unaligned with human preference and is often deemed unsafe [31, 48]. These unsafe models can be used for example to get instructions for building a bomb or other malicious activities. To improve instruction following and enhance safety, model creators perform an alignment fine-tuning stage. Usually, only the aligned version of the LLM is published, and the recovery of the original Pre-FT unsafe weights, is implicitly assumed to be impossible. While for existing models the recovery of the Pre-FT weights poses a security and safety vulnerability; for future superhuman models, it may lead to catastrophic consequences.
@@
Motivated by the above, we propose the task of Pre-Fine-Tuning Weight Recovery. In this paper, we tackle this task in cases where multiple LoRA fine-tuned flavors of the same source model are available. We present an overview of our setting in Fig. 1. This task is particularly timely due to two trends: i) Popular foundation models come in multiple flavors. E.g., LLaMA 2, Code LLaMA, Code LLaMA-Python, Code LLaMA-Instruct. ii) LoRA is becoming a key component for creating SoTA models [27, 46]. These two trends have not yet merged, i.e, we are not aware of multi-flavored foundational models that use LoRA alignment fine-tuning. Here, we bring to the attention of the community the risks and perils involved in merging these trends.
@@
Figure 2: Mistral DPO Results: Our method, Spectral DeTuning, recovers the pre-fine-tuning generation capabilities with high precision, essentially undoing the DPO alignment LoRA fine-tuning. In green exact recovery, in red unrecovered words. Best viewed in color
@@ 4 Problem Definition
Attack setting. The attacker has access to the weights of $n$ different $\mathcal{F}_{\mathcal{P}}$ models, all LoRA fine-tuned from the same pre-trained source model $\mathcal{P}$. The attack succeeds with precision $\epsilon$ if the attacker can accurately recover the weights of the pre-trained source model $\mathcal{P}$ up to an $\epsilon$ precision. Formally, given $\left\{\{W_{i}^{\prime(j)}\}_{j=1}^{m}\right\}_{i=1}^{n}$, the attacker needs find $\{W^{*(j)}\}_{j=1}^{m}$ such that
@@ 5 Spectral DeTuning
We now describe our method for carrying out a Pre-FT weight recovery attack. We start by introducing our optimization objective, followed by our optimization method and finally, a rank scheduler that stabilizes the optimization and results in better convergence. For simplicity, assume for now that all $n$ LoRA fine-tuned models used the same rank $r$, and that the value of $r$ is known to the attacker, in Secs. 5.3 and 5.4 we relax these assumptions. For brevity, we omit the layer index superscript $(j)$ and perform the same optimization across all layers independently.
@@ 5.2 Pre-FT Weight Recovery Algorithm
Where $\Sigma_{i|r}$ represents the top $r$ singular values of $\Sigma_{i}$.
W-step. By fixing the values of $M^{*}_{1},..,M^{*}_{n}$, we can easily compute the optimal value of $W$. The optimization problem is given by:
@@ 5.3 Rank Scheduler
The algorithm proposed in Sec. 5.2 tends to perform well in general. However, we empirically found that solving the optimization problem with high ranks can result in slow and inaccurate convergence. We therefore introduce a rank scheduler. The idea of the rank scheduler is to start by forcing $M_{i}$ to be of rank $r^{*}<r$, allowing Spectral DeTuning to focus on the most significant principal components first. $r^{*}$ is increased according to a schedule until finally $r^{*}=r$. Specifically, we use an “Increase on Plateau” type of scheduler where the rank is increased whenever the loss term from Eq. 3 plateaus. When not all LoRAs have the same rank, we assign a distinct rank scheduler to each LoRA. The rank scheduler requires knowing the LoRA rank; we show how to estimate it in Sec. 5.4. For more details see App. F. We show empirically in Sec. 8 that there are cases where the rank scheduler improves the rate and quality of convergence significantly.
@@
5.4 LoRA Rank Estimation
We propose an effective heuristic for estimating LoRA rank. Assume we have two LoRA fine-tuned models $W^{\prime}_{i}=W+M_{i}$ and $W^{\prime}_{j}=W+M_{j}$, where the ranks of $M_{i},M_{j}$ are $r_{i},r_{j}$ respectively. While it is not trivial to recover the rank of $M_{i}$ solely by observing $W^{\prime}_{i}$, there is a trick. Subtracting the two fine-tuned models obtains $W^{\prime}_{i}-W^{\prime}_{j}=M_{i}-M_{j}$. Importantly, the rank $W^{\prime}_{i}-W^{\prime}_{j}$ is upper bounded by $r_{i}+r_{j}$, i.e., $rank(W_{i}^{\prime}-W_{j}^{\prime})\leq r_{i}+r_{j}$. Given $n$ LoRAs, there are $\frac{n(n-1)}{2}$ distinct inequalities for the $n$ unknown ranks $r_{1},r_{2},..,r_{n}$.
We can formulate this as a linear programming problem as follows:
$\begin{array}[]{ll}\underset{\mathbf{r}}{\text{minimize}}&\mathbf{1}^{T}\mathbf{r}\\ \text{subject to}&\mathbf{A}\mathbf{r}\geq\mathbf{b}\\ &r_{i}\geq 1,\quad\forall i\end{array}$
where:
  • $\mathbf{A}\in\{0,1\}^{n^{2},n}$ so that $A_{ni+j,i}=1$ and $A_{ni+j,j}=1$ and 00 elsewhere.
  • $\mathbf{b}\in R^{n^{2}}$ so that $b_{ni+j}=rank(W^{\prime}_{i}-W^{\prime}_{j})$.
In practice, we populate $b$ using a numerical rank computed via the multiplicative gap following a similar protocol to [6]. Using an off-the-shelf linear programming solver accurately retrieves the ranks. We demonstrate the accuracy of this method in Sec. 8, the unknown ranks were recovered perfectly in all cases.
@@ 6.1 Dataset
For each source model, we curate $15$ LoRA models fine-tuned on diverse datasets, tasks, and objectives. The dataset comprises a diverse array of layer types, including self-attention, cross-attention, and MLPs. This diversity enables us to assess the generalization capabilities of Pre-FT methods. The evaluation can be conducted on a per-model basis, per layer type, or per layer depth, allowing for a comprehensive analysis of Pre-FT methods. Overall, our dataset includes $544$ source model layers. When taking into account the fine-tuned LoRA layers, the dataset includes over $8,000$ layers. For further details see App. E.
@@ 8 Ablations
Robustness to Unknown and Varying Ranks. We tested the LoRA rank estimation heuristic presented in Sec. 5.4 on hundreds of combinations of LoRAs with different ranks. The heuristic achieved an accuracy of $100\%$. We further tested the idea of using a dedicated rank scheduler for each LoRA model as described in Secs. 5.3 and 5.4. We use $n=6$ fine-tuned LoRAs with ranks $[8,32,32,32,64,100]$ taken from CivitAI. Spectral DeTuning is robust to the varying ranks, exhibiting only a minor decrease in performance despite the higher rank of the LoRAs (See Tab. 5).
Robustness to Different Models. We demonstrate the robustness of Spectral DeTuning to cases where a fine-tuned LoRA from a different Pre-FT model (with the same architecture) was mixed into the set of fine-tuned LoRAs. Using the same heuristic presented in Sec. 5.4, the difference between the mixed model weights and any other LoRA should be of full rank (since the Pre-FT model is different) and trivial to detect.
We validated this solution using Stable Diffusion. We added to the set of fine-tuned LoRA models a model that originated from Stable Diffusion 1.4 (all the others originated from Stable Diffusion 1.5). Indeed, the above steps indicated the LoRA that originated from Stable Diffusion 1.4 has a full rank difference from any other LoRA (while the pairwise rank between the LoRAs that used the same Pre-FT model were low rank, as expected). This allows us to detect the LoRA that got mixed up into the set and remove it.
@@
Table 5: Robustness to Unknown and Varying Ranks Results: We test the robustness to LoRAs with varying ranks. Spectral DeTuning is robust to varying ranks, exhibiting only a minor decrease in performance. We use $n=6$ fine-tuned LoRAs with ranks $[8,32,32,32,64,100]$ taken from an online LoRA marketplace
Method W-Error $\downarrow$ LPIPS $\downarrow$
LoRA FT -5.882 0.462
Mean LoRA -6.969 0.307
Spectral DeTuning -14.453 0.073
@@ 10 Conclusion
11 Acknowledgements
This work was supported in part by the “Israel Science Foundation” (ISF), the “Council for Higher Education” (Vatat), and the “Center for Interdisciplinary Data Science Research” (CIDR).
@@ Appendix A The Effect of the Number of LoRAs on Semantic Convergence
Appendix D LoRA Rank vs. W-Error
In Tabs. 7 and 7 we show the results for the ViT model when using different LoRA ranks and fixing the number of LoRAs.
Table 6: Using $5$ LoRAs
Rank W-Error
8 $-15.636$
12 $-15.550$
16 $-13.480$
32 $-4.817$
Table 7: Using $5$ LoRAs
Rank W-Error
8 $-15.822$
12 $-15.773$
16 $-15.258$
32 $-9.639$
Appendix E LoWRA Bench Dataset
@@ E.2 Mistral Models
Following the original setup, use a rank $r=64$ and LoRA fine-tune the q_proj, k_proj, v_proj, and o_proj layers. This protocol results in $128$ Pre-FT model layers and a total of $128\cdot 15=1920$ LoRA fine-tuned layers for both the SFT and DPO stages. See Tabs. 10 and 10 for the fine-tuning hyper-parameters. For inference we use the following decoding hyper-parameters: max_new_tokens=50, do_sample=True, temperature=0.7, top_k=50, top_p=0.95.
@@
E.3 Stable Diffusion Models
@@
Appendix F Spectral DeTuning Implementation Details
@@
Appendix G Runtime and Compute
@@
Appendix H Detecting the Fine-Tuned Layers
@@
Appendix I Algorithm with Rank Scheduler
@@
Appendix J Mistral Additional Results
@@
Appendix K Stable Diffusion Additional Results